# ChemBOMAS: Accelerated BO in Chemistry with LLM-Enhanced Multi-Agent System

### v1

@@ 5 Experiment
To rigorously assess the effectiveness and delineate the superiority of our approach, we conducted a comprehensive validation process. This involved extensive computational (dry lab) simulations utilizing multiple benchmarks, alongside a previously unreported wet laboratory experiment. The computational simulations proceeded in several stages. First, we conduct a direct comparison of our method against a traditional BO algorithm on a long-iteration chemical optimization task. Then, to assess our method’s robustness in mitigating the ‘cold start’ phenomenon inherent in BO, it was benchmarked against three representative BO techniques across four distinct benchmarks, with a specific focus on early-stage performance. Finally, an ablation study was conducted to examine the contributions of individual components of our proposed system to the overall optimization performance. For wet laboratory validation, our approach was applied to a complex chemical reaction optimization task designed by a pharmaceutical company. Its performance was then benchmarked against the results obtained by experienced human chemical experts.
5.1 Benchmark evaluations
5.1.1 Experimental Setup and Benchmarks
This subsection details the setup for the computational evaluations, including ChemBOMAS configurations and benchmark dataset characteristics, crucial for understanding the subsequent comparative analyses and ablation study. The configurations for the baseline algorithms (Traditional BO, LLAMBO Liu et al. [2024], LA-MCTS Wang et al. [2020]) are provided in Supplementary Material
For the ChemBOMAS framework, general LLM-driven tasks, such as knowledge extraction and reasoning, utilized the GPT-4O API and GPT-O1 API. A critical measure to prevent data leakage involved manually curating the literature database to ensure no publications directly corresponding to the benchmark datasets were included. Furthermore, prompts were meticulously engineered to compel the LLM to derive insights and predictions based solely on the provided literature corpus, thereby minimizing reliance on its pre-trained knowledge about the specific benchmark reactions.
The exploration constant () for the UCB algorithm in the coarse-grained optimization stage was set to 10, consistent with the LA-MCTS configuration. All other Bayesian Optimization parameters for the fine-grained stage were identical to those of the baseline method. The data module was constructed by fine-tuning an 8B parameter LLaMa 3.1 model. Our proposed method was evaluated using high-throughput screening reaction datasets curated from the literature. While these datasets are of moderate size, containing up to approximately 6,000 reactions, they possess a critical attribute for rigorous evaluation: completeness within the defined variable space. Specifically, for every combination of reaction parameters proposed by the optimization framework, a corresponding experimental result is available. This unique completeness enables a deterministic and quantitative assessment of our method’s recommendations. Furthermore, the integrity and consistency of these datasets are assured as all data were generated on automated experimental platforms, thereby minimizing human-induced error and experimental bias. The reliability of these benchmarks is further underscored by the adoption of several of these datasets in multiple prior studies on reaction optimization Shields et al. [2021], Wang et al. [2024b].
5.1.2 Benchmark Results
Long iteration As shown in figure 2 A, our method achieved the best objective value that exceeded the traditional BO baseline by 30% in the initial round and by 17% over the entire optimization campaign. Notably, ChemBOMAS reached convergence by the 4th experimental round, representing over a 50% improvement in optimization efficiency compared to the traditional BO, which had not yet converged by the 9th round. Moreover, Figure 2 B illustrates that ChemBOMAS consistently outperforms the baseline, as confirmed by ten independent repetitions of the optimization campaign.
Short iteration As shown in Figure S1 in supplementary material, our benchmark evaluations revealed three key findings. First, ChemBOMAS consistently outperformed all baselines in initial performance across the four chemical reaction optimization tasks. Second, it pinpointed the optimal objective value for each task within just four iterations. Third, in the Suzuki coupling task, LLAMBO reached an optimal target value of 96.3% in the first iteration, which was slightly higher than the initial ChemBOMAS value of 92.56%, possibly due to the abundant relevant literature that guides LLAMBO in selecting high-value candidates at the early stage. However, at the end of the optimization campaign, ChemBOMAS matched LLAMBO’s optimal objective value of 96.3%.
Ablation study The ablation study results, presented in Figure 2 C, demonstrate the superior optimization performance of the complete ChemBOMAS framework. The complete ChemBOMAS outperformed three distinct ablated versions: (i) without the Knowledge Module, (ii) without the Data Module, and (iii) without either module. A key observation is that the absence of the Data Module led to better initialization performance than the absence of the Knowledge Module. This suggests a more significant contribution of the Knowledge Module to alleviating the cold-start problem in BO. Nevertheless, the standard BO, devoid of the two specialized modules, exhibited the lowest optimization performance among all tested configurations.
5.2 Wet Experiments
5.2.1 Experiment Results
As shown in Figure 3, during the wet experiment task, ChemBOMAS successfully identified the optimal reaction condition with a yield of 96%, markedly outperforming the 15% yield achieved by a chemist employing the traditional control variable method. Additionally, three noteworthy phenomena emerged. First, in the initial round, ChemBOMAS had attained a maximum product yield of 90%, surpassing the target threshold of 75%. Second, the optimal reaction condition yielding 96% was discovered in the early stage of the optimization process, specifically in the second iteration. Third, as the optimization progressed, ChemBOMAS increasingly recommended reaction conditions with yields exceeding the 75% target threshold, indicating a continuous refinement of the surrogate model. The number of high-yielding conditions (75%) identified in rounds one through five was one, two, three, five, and five, respectively. The strong initialization performance, rapid convergence, and progressive model improvement collectively demonstrate the effectiveness and efficiency of ChemBOMAS in accelerating chemical reaction optimization. Additionally, to validate the robustness of ChemBOMAS’s initialization performance, the initial-round sampling was repeated ten times with the fixed experimental configurations, as detailed in Supplementary Materials. In the ten repeated initialization tests using ChemBOMAS, each run consistently identified at least two reaction conditions with yields exceeding 60%. Moreover, reaction conditions achieving yields above 80% appeared in 70% of the validation tests, totaling 11 such high-yield conditions across all trials. These results demonstrate that ChemBOMAS reliably mitigates the “cold-start” problem inherent to BO optimization.
@@ 6 Limitations
While ChemBOMAS demonstrates significant advancements in accelerating Bayesian Optimization for chemical reactions, several limitations warrant discussion and offer avenues for future research:
Dependence on LLM Performance and Knowledge Base: The efficacy of the knowledge-driven coarse-grained optimization heavily relies on the LLM’s ability to accurately interpret chemical literature, rank variable importance, and identify key physicochemical properties. Errors, biases in the LLM’s reasoning, or the quality of available chemical literature and databases for a specific reaction class could lead to suboptimal initial search space decomposition, potentially hindering the overall optimization efficiency.
Scalability to Extremely High-Dimensional Spaces: The current wet-lab validation involved a six-dimensional parameter space. While this is significant, some chemical optimization problems, particularly in materials discovery or complex biological systems, can involve even higher dimensionality. The effectiveness of the UCB-guided tree search and the computational cost of LLM inference for pseudo-data generation across all subspaces might face challenges in such scenarios, although the paper notes that LLM inference is negligible compared to wet-lab experiments.
@@ 7 Conclusion
ChemBOMAS presents an LLM-enhanced multi-agent framework designed to accelerate Bayesian Optimization in the context of chemical reactions. Through a synergistic combination of knowledge-driven coarse-grained search space decomposition and data-driven fine-grained optimization, this approach seeks to mitigate common challenges like data scarcity and complex reaction mechanisms. Results from benchmark evaluations, along with encouraging outcomes from wet-lab validation on a demanding, previously unreported industrial reaction—where ChemBOMAS showed improved performance compared to domain expert methods—suggest its potential for practical application. ChemBOMAS offers a promising direction for facilitating chemical discovery and enhancing the optimization of complex chemical processes.
@@ Appendix C Benchmark Detail
This section provides further details on the benchmark datasets used for evaluating ChemBOMAS. These datasets represent common chemical reaction optimization tasks and have been previously utilized in the literature for benchmarking machine learning-driven optimization approaches. The key characteristics of these datasets are summarized in Table 1-3.
| Dataset Name | Total Number | Reaction variables | Objective |
| Suzuki coupling Perera et al. [2018] | 5760 | 4 (Reactants, Ligand, Base, Solvent) | yield |
| Buchwald-Hartwig coupling Ahneman et al. [2018] | 3955 | 4 (Ligand, Base, Additive, Solvent) | yield |
| Tandem arylation Qiu et al. [2022] | 1032 | 3 (Reactants, Solvent) | yield |
| Enantioselectivity(Catalyst, Reactants,) Zahrt et al. [2019] | 1076 | 3 (Reactants, Base) | selectivity |
Appendix D Wet Experiment Detail Protocol
| Rules | |
| Catalyst Pd Usage | 0.002 mmol |
| Ligand P Usage | 0.008 mmol |
| Base Usage | 0.3 mmol |
| Reactant 1 Usage | 0.2 mmol |
| Reactant 2 Usage | 0.25 mmol |
| Solvent Usage | 2 mL |
| Reaction Time | 24 h |
| Variable Dimensionality | Variable Number | Subset Number | Division Basis |
| Catalyst Type | 5 | 2 | Valence |
| Ligand Type | 14 | 6 | Steric and Electronic Effect |
| Base Type | 14 | 3 | Basicity |
| Solvent Type | 11 | 3 | Polarity |
| Water usage | 3 | 3 | Solubility |
| Temperature | 3 | 3 | Activation Energy |

### v2

@@ 4 Experiment
4.1 Data
The pre-training phase employed a subset of the Pistachio dataset containing approximately 50,000 chemical reaction entries, none of which contained objective performance labels.
For the fine-tuning and Bayesian Optimization (BO) phases, three benchmark datasets were used: Suzuki perera2018platform/suzuki_dataset, Arylation shields2021bayesian/arylation_dataset, and Buchwald ahneman2018predicting/buchwald_dataset. In each case, only 1% of the labeled data was randomly selected for fine-tuning the LLM regressor; the effectiveness of this data volume is analyzed in Appendix I.2. For the Buchwald dataset, which exhibits inconsistencies in reactants and products across entries, two consistent subsets were constructed, denoted Buchwald and Buchwald, to serve as rational benchmarks for the optimization task. Detailed statistics for all benchmarks are provided in Appendix C.
4.2 Experiment Setup
The LLM regressor in data-driven module was trained on NVIDIA A800 GPUs. For fine-tuning the LLM regressor, the hyperparameters were set as follows: learning rate of , batch size of 24, and 100 training epochs.
In the knowledge-driven module, the search tree was constructed using a UCB policy with an exploration constant . The BO process was run for 40 iterations, initialized with 1% of the data as the prior and sampling 0.1% of the dataset in each iteration. The acquisition function and other BO configurations were kept consistent with the baseline methods for a fair comparison. Each optimization experiment was independently repeated 5 times with different random seeds, and the average performance across these runs is reported as the final result. (The prompts for LLM clustering and further implementation details are provided in the Appendix G.)
4.3 Performance Comparsion
4.3.1 Regression Models
The quality of the pseudo-data is directly influenced by the prediction accuracy of the regression model. We evaluated the performance prediction accuracy of ChemBOMAS against three categories of existing regression models on three chemical datasets: 1) general-purpose LLMs (GPT series) with zero-shot inference; 2) open LLMs fine-tuned on 1% labeled data; and 3) scientific LLMs with molecule pre-training fine-tuned on 1% labeled data . The prediction metrics for each model are summarized in Table 1, from which two key observations can be drawn.
Model Suzuki Arylation Buchwald MSE MAE R MSE MAE R MSE MAE R General-purpose LLMs with zero-shot inference GPT-4o 2207.17 40.02 -1.80 2702.58 44.86 -2.63 1512.44 33.60 -1.03 GPT-5 1218.93 30.34 -0.55 1515.81 33.68 -1.04 1516.62 33.55 -1.04 Open LLMs fine-tuned on 1% labeled data Qwen3-7B 820.48 22.10 -0.04 848.51 22.52 -0.14 998.22 25.25 -0.34 GLM4-9B 593.49 18.94 0.25 739.20 20.78 0.01 719.72 20.77 0.00 LLaMa-3.1-8B 685.55 20.50 0.13 679.72 19.57 0.09 739.27 20.57 0.01 Scientific LLMs with molecule pre-training and fine-tuned on 1% labeled data MolT5-Large 1081.23 25.13 -0.37 1094.86 25.40 -0.47 1098.16 25.37 -0.47 Galactica-1.3B 727.18 22.23 0.08 785.01 21.83 -0.05 857.79 22.54 -0.15 ChemBOMAS 633.68 19.47 0.20 650.00 19.55 0.13 593.76 18.52 0.20
First, ChemBOMAS demonstrated superior effectiveness and versatility in chemical performance regression. As shown in Table 1, ChemBOMAS achieved the highest prediction accuracy on the Arylation and Buchwald datasets, with scores exceeding the second-best model by 2000% and 140%, respectively. On the Suzuki dataset, ChemBOMAS outperformed six of the seven compared models, ranking second only to GLM4-9B. However, the poor generality of GLM4-9B is evident from its near-zero scores on the other two datasets.
Second, task-specific fine-tuning proved essential. Despite their general capabilities, the off-the-shelf general-purpose LLMs GPT-4o openai2024gpt4ocard and GPT-5 OpenAI2025GPT5 performed poorly on this specialized regression task, consistently yielding strongly negative scores, lower than most fine-tuned models, which also confirms that these chemical datasets were not part of their training data. Among the open-source models, LLaMa-3.1-8B grattafiori2024llama3herdmodels exhibited a favorable balance of prediction accuracy and generalization, justifying its selection as the base model for the data-driven module of ChemBOMAS.
Furthermore, we investigated the impact of fine-tuning data volume on pseudo-data quality and BO performance (see Tables 7 and 8). The predictive performance improved gradually as data volume increased from 0.25% to 4%. Notably, the value first turned positive and consistently exceeded 0.1 across all datasets at the 1% volume. Table 8 indicates that BO’s optimization performance does not linearly correlate with the regression model’s ; a value between 0.1 and 0.2 is sufficient for BO to identify high-performing conditions. Therefore, using 1% data volume for fine-tuning represents a rational and effective trade-off between cost and performance.
4.3.2 Cluster Methods
We evaluated the impact of the search tree structure on BO by comparing scenarios with and without a tree, as well as trees constructed using three distinct strategies: expert guidance (Expert Guided), data-driven approach (ChemBOMAS), and knowledge-driven approach (ChemBOMAS). All methods were initialized with an identical, fixed set of pseudo-data to ensure a fair comparison.
Suzuki Arylation Method Best Found (%) Initial (%) 95% Max Iter Best Iter Best Found (%) Initial (%) 95% Max Iter Best Iter BO 83.16 54.04 16 38 80.45 64.64 8 35 Expert-Guided 96.15 61.96 3 3 79.67 78.71 1 36 ChemBOMAS 89.51 72.98 29 29 79.67 67.75 2 28 ChemBOMAS 96.15 72.98 3 3 82.23 78.71 1 39 Buchwald Buchwald Method Best Found (%) Initial (%) 95% Max Iter Best Iter Best Found (%) Initial (%) 95% Max Iter Best Iter BO 78.86 45.39 26 31 56.29 27.96 13 34 Expert-Guided 79.71 75.55 2 32 56.81 53.33 2 2 ChemBOMAS 79.68 79.04 1 6 56.81 56.81 1 1 ChemBOMAS 79.77 75.55 2 11 56.81 56.81 1 1
The results, summarized in Table 2, demonstrate that employing a subspace tree search significantly accelerates Bayesian Optimization, yielding up to a 34-fold improvement in optimization efficiency compared to the standard BO baseline without a tree. Notably, the clustering methods derived from ChemBOMAS—both ChemBOMAS and ChemBOMAS—consistently matched or surpassed the performance of the expert-guided approach across all benchmarks. This underscores the robustness and reliability of our automated framework for variable space decomposition. Furthermore, the knowledge-driven clustering strategy achieved superior optimization performance on more benchmarks compared to its data-driven counterpart, highlighting the value of incorporating structured chemical knowledge.
4.3.3 Optimization
As shown in Figure 2, ChemBOMAS demonstrates consistent and superior performance over all baseline methods across the four benchmark datasets in terms of optimal result, convergence rate, initialization performance, and robustness.
In terms of final performance and convergence speed, ChemBOMAS identified the highest objective values—96.15% (Suzuki), 81.26% (Arylation), 80.00% (Buchwald) and 56.80% (Buchwald)—achieving convergence in just 3, 39, 23, and 2 iterations, respectively. This represents the fastest convergence among all methods.
Regarding initialization performance, ChemBOMAS attained the highest initial performance on the Arylation, Buchwald, and Buchwald datasets. Although it started 5.12% lower than LA-MCTS and BO-ICL on the Suzuki dataset, it surpassed all baselines by the third iteration and proceeded to converge, highlighting its strong optimization capability.
Two additional observations further underscore the robustness of ChemBOMAS. First, it exhibited the lowest variance across five independent optimization runs, indicating high stability. Second, its performance remained consistently effective and was largely unaffected by the sampling batch size (see Appendix I.1 for details). These findings collectively confirm the reliability of our method.
To evaluate the generality of ChemBOMAS beyond chemistry, we assessed its optimization performance on a materials science benchmark. As shown in Table 9 (see Appendix J for dataset details), ChemBOMAS maintains competitive performance, demonstrating its applicability to other scientific domains.
To further validate the practical applicability of ChemBOMAS and preclude the possibility of knowledge leakage in the LLM, we conducted wet-lab optimization for a previously unreported chemical reaction (see Appendix H for details). As shown in Figure 5, ChemBOMAS identified the optimal reaction condition with a yield of 96% after evaluating only 43 samples in 2 iterations. This result markedly outperforms the 15% yield obtained by a chemist using the traditional control variable method, demonstrating the framework’s effectiveness in real-world scenarios.
4.4 Ablation Study
We first evaluate the impact of pre-training and fine-tuning on the regression performance of ChemBOMAS. The prediction accuracy is measured on the Suzuki, Arylation, and Buchwald datasets using Mean Squared Error (MSE), Mean Absolute Error (MAE), and the coefficient of determination ().
Model Configuration Suzuki Arylation Buchwald MSE MAE MSE MAE MSE MAE w/o Pre & SFT 2338.02 39.43 -1.966 1797.88 32.54 -1.413 1881.96 33.73 -1.527 Pre-train Only (w/o SFT) 2407.22 40.27 -2.054 1853.70 33.24 -1.488 1795.72 32.52 -1.411 SFT Only (w/o Pre-train) 685.55 20.50 0.130 679.72 19.57 0.088 739.27 20.57 0.007 Pre-train & SFT 633.68 19.47 0.196 650.00 19.55 0.128 667.16 19.51 0.104
The results in Table 3 indicate that the combined use of pre-training and supervised fine-tuning (SFT) yields the best predictive performance across all benchmarks. Notably, SFT alone (without pre-training) achieves the second-best performance and substantially outperforms models using only pre-training or those without any training. This strongly suggests that supervised fine-tuning is the most critical component for adapting large models to chemical performance prediction tasks.
We further evaluate the contribution of each module by comparing the complete ChemBOMAS framework against three ablated versions: (i) without the data-driven module, (ii) without the knowledge-driven module, and (iii) without both modules. The results in Table 4 demonstrate that both modules are critical to the framework’s performance. Ablating either module leads to a significant degradation in both optimization efficiency and final effectiveness.
For instance, on the Suzuki dataset, the full ChemBOMAS achieves the optimal value of 96.15% within only three iterations. In contrast, removing the data-driven module reduces the optimum to 82.65%, while disabling the knowledge-driven module limits it to 88.98%. The performance of the single-module ablations is comparable to or only marginally better than the version lacking both modules, indicating that neither strategy alone is sufficient. These results underscore that the synergy between the knowledge-driven and data-driven strategies is essential for creating a highly efficient and robust optimization framework.
Suzuki Arylation Method Best Found (%) Initial (%) 95% Max Iter Best Iter Best Found (%) Initial (%) 95% Max Iter Best Iter Full ChemBOMAS 96.15 72.98 3 3 81.26 78.71 1 39 w/o data module 82.65 60.44 13 29 80.65 45.40 13 40 w/o knowledge module 88.98 65.95 21 37 79.67 45.98 10 40 w/o both 83.16 54.04 16 38 80.45 64.64 8 35
Buchwald Buchwald Method Best Found (%) Initial (%) 95% Max Iter Best Iter Best Found (%) Initial (%) 95% Max Iter Best Iter Full ChemBOMAS 80.00 75.55 2 23 56.81 53.33 2 2 w/o data module 79.90 59.32 10 40 55.53 33.12 17 32 w/o knowledge module 79.63 54.07 9 31 56.81 12.87 9 11 w/o both 78.86 45.39 26 31 56.29 27.96 13 34
@@ 5 Limitations
While ChemBOMAS represents a significant advancement in accelerating Bayesian Optimization for chemical reactions, its performance remains constrained by several factors. Most notably, the framework is inherently dependent on the accuracy and scope of the underlying LLM and its knowledge base; inference errors in literature parsing or incomplete corpora can lead to suboptimal search-space decompositions. In addition, the absence of explicit safety and feasibility constraints raises the risk of recommending theoretically optimal yet practically hazardous or infeasible conditions, underscoring the need for expert oversight or integration of safety-aware modules in future implementations.
@@ 6 Conclusion
ChemBOMAS presents an LLM-enhanced multi-agent framework designed to accelerate Bayesian Optimization in the context of chemical reactions. Through a synergistic combination of knowledge-driven search space decomposition and data-driven pseudo-data generation, this approach seeks to mitigate common challenges like data scarcity and complex reaction mechanisms. Results from benchmark evaluations, along with encouraging outcomes from wet-lab validation on a demanding, previously unreported industrial reaction—where ChemBOMAS showed improved performance compared to domain expert methods—suggest its potential for practical application. ChemBOMAS offers a promising direction for facilitating chemical discovery and enhancing the optimization of complex chemical processes.
@@ Appendix C Benchmark Detail
This section provides further details on the benchmark datasets used for evaluating ChemBOMAS.
C.1 Dataset for LLM Pre-train
The Pistachio dataset employed during the pre-training phase is a large-scale reaction information repository. Its core data was systematically extracted from the full texts of US patents (USPTO) and European patents (EPO) through automated text mining techniques. To enhance data diversity and accuracy, the dataset integrates information from multiple sources, including: - Structured data parsed from ChemDraw (CDX) files embedded directly within patent documents - Records sourced from specialized chemical databases such as Reaxys - Exported data from select electronic laboratory notebooks (ELNs) The dataset contains a total of 19.17 million chemical reactions. In this project, we primarily utilize the reaction SMILES strings for model pre-training.
C.2 Dataset for LLM Fine-tune
To conduct a rigorous and unbiased evaluation of model performance, we selected a series of publicly available benchmark datasets widely used in the field of chemical reaction optimization. The core strength of these datasets lies in their completeness: all were generated via high-throughput automated experimental platforms and encompass experimental results for every variable combination within a clearly defined chemical space (full factorial design). This exhaustive coverage effectively eliminates sampling bias, enabling deterministic quantitative evaluation of algorithmic recommendation performance against known experimental ground truth.
Specifically, we employed three recognized benchmark datasets: Suzuki, Arylation, and Buchwald reactions. During fine-tuning, we randomly sampled 1% of data from each dataset as training samples to adjust the pre-trained model.
Suzuki originates from the automated nanomolar-scale flow screening study reported by Perera et al. in 2018. The chemical space of the experiments comprised a full factorial combination of 4 halogenated quinolines, 3 boronic acid derivatives, 11 phosphine ligands, 7 bases, and 4 solvents. All reactions were conducted under uniform conditions (100 °C, 1-minute residence time, 9:1 organic/aqueous phase). Reaction yields were detected via dual UPLC-MS online detection and uniformly calibrated. The data is comprehensive and highly consistent, making it one of the widely adopted validation standards in the field.
Arylation was reported by Shields et al. in 2021 for Bayesian optimization studies. Its chemical space was generated via a full factorial design comprising 12 phosphine ligands, 4 bases, 4 solvents, 3 temperature gradients, and 3 concentration gradients. All experiments were conducted at high throughput in 96-well plates, with yields precisely quantified via UHPLC-MS coupled with internal standard methods. This dataset features no duplicates or missing data, exhibits uniform variable distribution, and has been validated by 50 practicing chemists, establishing it as a critical benchmark for optimizing C-H functionalization reactions.
Buchwald was published by Ahneman et al. in 2018, this dataset aims to predict yields of C-N coupling reactions via machine learning. Experiments were conducted in nanomolar-scale high-throughput format using 1536-well plates, systematically examining all combinations of 15 aryl halides, 4 ligands, 3 bases, and 23 isoxazole additives. All reactions proceeded under standard conditions (60 °C, DMSO, 16 hours), with yields quantified by LC-MS. This dataset is complete with no missing values, serving as an authoritative open-access resource for studying additive effects and modeling complex reaction systems.
C.3 Dataset for Bayesian Optimization
| Statistic | Suzuki | Arylation | Buchwald | Buchwald |
| Total data points (N) | 5030 | 3678 | 629 | 765 |
| Maximum Yield (%) | 96.15 | 84.65 | 80.91 | 56.81 |
| Minimum Yield (%) | 0.00 | 0.00 | 0.00 | 0.00 |
| Mean (%) | 33.04 | 29.05 | 42.24 | 18.71 |
| Median (%) | 26.86 | 25.53 | 42.21 | 11.34 |
| Standard Deviation (%) | 22.47 | 23.79 | 22.86 | 18.98 |
| 25% Quantile (%) | 15.26 | 6.87 | 23.14 | 0.72 |
| 75% Quantile (%) | 51.27 | 47.14 | 63.01 | 38.77 |
For Bayesian optimization tasks, we employed four benchmark datasets: Suzuki, Arylation, Buchwald, and Buchwald. The latter two originate from partitions of the aforementioned Buchwald-Hartwig dataset. To ensure consistency of target products within the optimization space, the original dataset was first divided into five independent subsets based on product molecular structures. We observed distinct high-yield and low-yield patterns in the reaction yields of these subsets. To ensure comprehensive evaluation, we selected one representative subset from each category, naming them Buchwald and Buchwald, respectively. Table 5 summarizes key descriptive statistics for these four datasets, while Figure 3 visually depicts their respective yield distributions via kernel density estimation (KDE) plots. These datasets exhibit distinct statistical characteristics, with average yields ranging from 18.71% to 42.24% and diverse distribution shapes. Collectively, they form a challenging optimization problem that effectively tests algorithm performance across varying data environments.
C.4 Comparative Algorithms
To evaluate the efficacy of our proposed method, we benchmark it against four algorithms that represent diverse approaches to black-box optimization.
Traditional Bayesian Optimization (BO) is a sequential optimization method that utilizes a surrogate model to approximate the objective function and an acquisition function to guide subsequent sampling. The baseline implementation in this work employs a Gaussian Process (GP) with a Matérn kernel as the surrogate, a qLogEI acquisition function, and one-hot encoding for inputs. BO did not adopt more complex encoding schemes because prior research indicated their benefits were negligible Taylor23, Shields2021.
Bayesian Optimization with In-Context Learning (BO-ICL) is a method that integrates a frozen Large Language Model (LLM) with BO, as proposed by ramos2023bayesian/boicl. It leverages the LLM’s in-context learning ability to form a surrogate model by translating experimental parameters into textual prompts, thereby predicting outcomes and uncertainty without model fine-tuning.
Gaussian Process Optimized LLMs (GOLLuM) is a method that achieves a deeper fusion of LLMs and Bayesian optimization, introduced by rankovic2025gollum. It utilizes the LLM’s embedding space as a deep kernel for a GP and jointly optimizes the embedding and GP hyperparameters to learn a task-specific representation space.
Latent Action Monte Carlo Tree Search (LA-MCTS) is a meta-algorithm for high-dimensional optimization, developed by wang2020learning/lamcts. It employs Monte Carlo Tree Search to dynamically partition the search space into high- and low-performance regions and subsequently deploys a local optimizer within promising subregions.
Appendix D Dual-Strategy Refinement for Enhanced Optimization
To mitigate the detrimental influence of noise and redundancy inherent in generated pseudo-data, we introduced a dual-pronged refinement strategy. This approach was designed to dynamically curate the pseudo-dataset, ensuring its quality and diversity throughout the optimization process. The strategy combined a local, similarity-based removal mechanism with a global, performance-driven pruning policy. This ensured that the pseudo-dataset remained a reliable and informative asset for guiding the optimization, particularly in complex search spaces.
Data Similarity (Local Removal): We utilized the final token embedding, , to calculate cosine similarity. Upon acquiring a new real data point , the pseudo-dataset was updated by removing points that were too similar:
where was a predefined similarity threshold.
Observed Performance (Global Removal): As the optimization progresses, the model should be encouraged to explore more broadly. Therefore, based on the predicted performance values of the pseudo-points, we randomly discarded a proportion of pseudo-data, starting from those with high predicted performance downwards. The probability of discarding a pseudo-point was a monotonically increasing function of its predicted performance.
These generated pseudo-points could also provide further support for the construction of the knowledge-guided optimization tree in Stage 1. By adjusting the LLM’s temperature parameter during generation, we could produce a set of candidate tree structures, . Using the pseudo-points, we quantitatively evaluated these candidates. Let be the set of leaf nodes of tree , and let be the subset of pseudo-points belonging to node . The tree structure that minimized the weighted average of intra-node variances was selected as the optimal one:
This ensured the selection of a tree that best partitions the search space into regions of homogeneous performance, guiding the subsequent optimization more effectively.
Appendix E Qualitative Comparison of Optimization Trajectories
To qualitatively assess how well the automated clustering strategies of ChemBOMAS emulate expert-level reasoning, we visualized the optimization progress. Figure 4 provides a comparative heatmap of the "Best Found" objective value over 40 iterations for three different search tree configurations: one guided by human experts, one by our knowledge-driven module (ChemBOMAS), and one by our data-driven module (ChemBOMAS).
The visual evidence strongly suggests that both automated ChemBOMAS strategies produce optimization trajectories that are remarkably consistent with the expert-guided approach. The color progression—from blue (lower values) to red (higher values)—is highly similar across all three methods. This indicates that the subspaces identified as promising by the LLM-driven modules align well with those selected by human domain experts. The ability of both the knowledge-driven and data-driven variants to rapidly progress towards high-yield regions in a manner analogous to the expert baseline underscores the effectiveness of our framework in automatically structuring the search space in a chemically meaningful way. This qualitative alignment provides further confidence in the robustness and practical utility of ChemBOMAS for real-world chemical optimization tasks.
Appendix F Complete Algorithm Process
To provide a comprehensive and formal description of the ChemBOMAS framework, we present its complete algorithmic process in Algorithm 1. This pseudocode encapsulates the synergistic, two-stage optimization strategy detailed in Section 3.
The algorithm begins by initializing a hierarchical search tree using the LLM-guided knowledge-driven approach as shown in Section 3.4. The main loop then iterates through the coarse-grained optimization phase, where a UCB policy navigates the tree to select a promising subspace. Within this selected subspace, the algorithm transitions to the fine-grained, data-driven optimization phase. Here, a standard Bayesian Optimization procedure is executed, but it is significantly accelerated by an informative prior constructed from both real experimental data and pseudo-data generated by the fine-tuned LLM regressor (Section 3.3).
After each experimental evaluation, the results are backpropagated to update the value estimates of the nodes in the search tree, refining the knowledge-driven search for subsequent iterations. This process continues until the predefined budget of evaluations is exhausted, ultimately returning the best-performing experimental configuration found.
Appendix G Details of Prompts
As outlined in the main text, our methodology leverages LLMs to support several critical tasks in reaction optimization, such as analyzing literature, assessing parameter significance, and understanding physicochemical properties to inform the construction of a hierarchical optimization tree. This appendix section presents a detailed overview of the specific prompts designed to guide the LLM in executing these crucial Tasks.
G.1 Prompt of Data Module
As detailed in the Section 3.3, our pre-training phase employs a conditional prediction task. Given the reactants and products, the model’s objective is to predict the corresponding reaction conditions. This process utilizes a Causal Language Modeling (CLM) loss, where the model learns to predict the next token in the sequence of reaction conditions.
To provide concrete examples of the input format for this task, this appendix section presents a selection of prompts utilized during the pre-training phase. These prompts typically consist of the reactants, products, and the target reaction condition sequence that the model is trained to predict. Furthermore, in line with the methodology described in the main text, these input sequences are augmented with functional group annotations (generated via RDKit) to enhance the model’s chemical awareness; the augmentation of the prompt is also reflected in the examples provided below.
Prompt of Condition Prediction Pretraining:
To further clarify the input structure for this prediction task, the following examples demonstrate the format used:
Prompt of Yield Prediction Fine-tuning:
To fine-tune LLM for precise prediction of chemical reaction yields, we combine key chemical information—including reaction type, products, reactants, and reaction conditions—into structured prompts. This approach guides the model to learn the complex relationships between these variables and reaction outcomes, enabling it to output a specific numerical prediction.
Below is an example prompt for yield prediction fine-tuning from the reactants in the Suzuki coupling dataset.
G.2 Prompts of Knowledge Module
The Knowledge Module, as described in Section 3.4, employs the LLM to systematically analyze chemical literature and physicochemical data. This involves ranking the impact of various reaction parameters and classifying components based on their physicochemical properties.
Variable Candidates Clustering Prompt: The prompt guides the LLM to identify key physicochemical properties of each variable and cluster variable candidates based on their similarity in the physicochemical properties. Below is an example of the prompt for variable candidates classification.
Appendix H Wet Experiments
H.1 Experiment Results
As shown in Figure 5, during the wet experiment task, ChemBOMAS successfully identified the optimal reaction condition with a yield of 96%, markedly outperforming the 15% yield achieved by a chemist employing the traditional control variable method. Additionally, three noteworthy phenomena emerged. First, in the initial round, ChemBOMAS had attained a maximum product yield of 90%, surpassing the target threshold of 75%. Second, the optimal reaction condition yielding 96% was discovered in the early stage of the optimization process, specifically in the second iteration. Third, as the optimization progressed, ChemBOMAS increasingly recommended reaction conditions with yields exceeding the 75% target threshold, indicating a continuous refinement of the surrogate model. The number of high-yielding conditions (75%) identified in rounds one through five was one, two, three, five, and five, respectively. The strong initialization performance, rapid convergence, and progressive model improvement collectively demonstrate the effectiveness and efficiency of ChemBOMAS in accelerating chemical reaction optimization.
H.2 Wet Experiment Detail Protocol
Additionally, to validate the robustness of ChemBOMAS’s initialization performance, the initial-round sampling was repeated ten times with the fixed experimental configurations, as detailed in Supplementary Materials. In the ten repeated initialization tests using ChemBOMAS, each run consistently identified at least two reaction conditions with yields exceeding 60%. Moreover, reaction conditions achieving yields above 80% appeared in 70% of the validation tests, totaling 11 such high-yield conditions across all trials. These results demonstrate that ChemBOMAS reliably mitigates the “cold-start” problem inherent to BO optimization.